RC Meeting and Run Plan, Friday, December 6th 2019

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links


  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in or comes out of laser-controlled access.
  • Make sure beam-size is between 100-500um sigma.
  • Still many TO day/swing shifts available! You don't need TO training, and will be guided on-site by the target experts!
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.

Experiment Status

Ongoing Issues

  • Helium leak in Moller warm return, (Dave Gaskell) [1]
    • Warmup completed and hose replaced, will do full cooldown on Monday [2]
    • Maybe ready for spin-dance on Tuesday
    •  ?PD asked: can we "guarantee" to participate if spin-dance on Wednesday?
  • Timing cuts and pre-run calibration (Eric, ongoing)
  • HMS/SHMS rotation:
    • rotated to 20 degrees and parked, will need access to go to 30 degrees
    • rotation can be certified only after work on pivot completed (next week)
  • Target Status
  •  ?Q: beam orbit lock procedure?
  •  ?Target names in target GUI and MCC consistent?
    • Can we use "carbon foil", "holy carbon" and "optics target" or do we need to be more specific when communicating with MCC?
  • Needed for startup/beam authorization
    • Beamline checklist (Gaskell, DONE)
    • Tech checklist (Walter, in the loop)
    • Detector checklist (Brad)
    • Target checklist (JP)
    • ERR signature from Patricia and Ed
    • Radcon checklist

Resolved Issues

  • Raster:
    • Min raster current for 1-2 pass set to 0, so don't need FSD mask.
    • Will use only single set of raster coils for 1-2 pass running to mitigate coil sync issue
  • Operational restrictions on targets set as decided in yesterday [3]
    • Verified with experts that Be+Al windows are setup correctly (as required by safety and used for previous experiments).
  • HMS/SHMS Magnets
    • Magnet control issues resolved [4]
    • Powered-on magnets to 2GeV, found issue with HMS Dipole NMR [5], Steve L. rebooted ioc (fixed).
    • Checked magnets today (including HMS dipole NMR), they are good to go [6]
  • Approximate Unser calibration available, good to go for non-intrusive BCM scan [7]
  • Timing cuts and pre-run calibration (Eric)

Run Plan and Scheduling

  • Target work going on during day today (JP/Todd/Simon/Ethan) to setup controls and finish alignment
  • Finish target work tonight
  • If good-to-go, inform Walter tomorrow at 7:00am that we need techs to help close up
  • Should be ready for beam by SWING Saturday at 2-pass
  • If target not ready/no beam authorization, try for Sunday, else Monday
  • Saturday DAY shift workers can assist with final hall work if needed
  • Participate in Spin Dance if possible (5-pass)

Startup run plan through weekend/Monday

  • Establish beam (~1-2 days) (empty target and carbon-hole)
  • Commission spectrometers (C-hole and/or 7-foil optics target)
  • Do 2-pass optics (1-2 shifts) (7-fold optics target)
  • Commission target (PolHe3 target)
  • Tentative 2-pass run plan [8]


Special Requests

  • Please be gentle with our target :-)